OpenStudy (anonymous):

The perimeter of a rectangle is 26cm one side is 9cm longer than the other side. Find the length of the sides.

OpenStudy (anonymous):

2(x + 9) + 2(x) = 26 x + x + 9 + 9 = x + x = 26 4x + 18 = 26 4x = 8 x = 2 So one side is 2 cm long, and the other side is 11 cm long.